OpenStudy (anonymous):

derivative of arcsin((5x+12(sqrt(1-x^2)))/13)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\frac{d}{dx} \sin^{-1} u = \frac{1}{\sqrt{1-u^2}} \frac{du}{dx}\]

OpenStudy (anonymous):

lae x=sin theta 5x+12sqrt(1-x^2)/13=sin(theta+arctan (12/5) so ans arctan 12/5

OpenStudy (anonymous):

its jsut a really messy chain rule

OpenStudy (anonymous):

lae x=sin theta 5x+12sqrt(1-x^2)/13=sin(theta+arctan (12/5) so ans is derivative of theta+ arctan 12/5= sin^-1x+c 1/sqrt(1-x^2)

OpenStudy (anonymous):

see if u sub. this x=sin theta then it is reduced greatly...

OpenStudy (anonymous):

^ I dont have a cule what u did at all lols

OpenStudy (anonymous):

clue*

OpenStudy (anonymous):

I am very sure you have something wrong

OpenStudy (anonymous):

\[\sin^{-1} (\frac{5x+12\sqrt{1-x^2}}{13})\]let \[x=\sin \theta\]\[\sqrt{1-x^2}\sqrt{(1-\sin^2 \theta)}=\cos \theta\]\[\frac{5x+12\sqrt{1-x^2}}{13}=\frac{(5\sin \theta+12 \cos \theta)}{13}=\sin(\theta +\tan^{-1} (\frac{12}5)\]\[\sin^{-1}(\sin (\theta+\tan^{-1}(\frac{12}5)))=\theta +\tan^{-1}(\frac{12}5)\] \[\frac d{dx}(\sin^{-1}x+\tan^{-1}(12/5))=\frac{1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

lol dont worry , your working still makes no sense , and you kind off got sided tracked and didnt answer the actual at all

OpenStudy (anonymous):

for example , the third last line of above post

OpenStudy (anonymous):

ohh I think i get third line

OpenStudy (anonymous):

but still you are doing too many steps in one line lols

OpenStudy (anonymous):

mr elecengineer if havin so much prob frm his workin y dnt u get ans fr me

OpenStudy (anonymous):

\[\frac{5\sin \theta +12 \cos \theta }{13}\]let \[\frac5{13}=\cos \alpha , \]so\[\frac{12}{13}=\sin \alpha\]\[\frac5{13}\sin \theta + \frac{12}{13}\cos \theta =\sin \theta \cos \alpha +\cos \theta \sin \alpha=\sin (\theta +\alpha)\]now \[\alpha=\tan^{-1}\frac{12}{13}\]

OpenStudy (anonymous):

elecengineer..............don't be so fast....!!!!!!!! it causes accident.!!!!!!!!!!!!

OpenStudy (anonymous):

I dont have any specific problem with it , just that I checked wolframa and the answer they give their is massive (obviously the result of chain rules )

OpenStudy (anonymous):

there is no short cut, when you change the variables by substitution then you must use even more chain rules when differentiating

OpenStudy (anonymous):

then it becomes so many chain rules that people start making mistakes

OpenStudy (anonymous):

look elec...i substituted only to reduce the given problem to a simplified form....

OpenStudy (anonymous):

has vaishali understood? thats the big query..

OpenStudy (anonymous):

its the fault of that person who makes mistake using chain rule ...not me...

OpenStudy (anonymous):

but I dont even get the whole point of your last line\[\frac{d}{dx} (\sin^{-1}(x) + \tan^{-1}( \frac{12}{5}) ) = \frac{1}{\sqrt{1-x^2}} \]

OpenStudy (anonymous):

hey guyz stop it

OpenStudy (anonymous):

like, isnt that obvious?, and it didnt answer the question

OpenStudy (anonymous):

I think you forgot what the actual question was lol

OpenStudy (anonymous):

elecengr is right..its actually simplified...theres no needo use chain rule

OpenStudy (anonymous):

if u sub x =sint it simply reduces to arcsin(sin(t + alpha)) which is t + alpha or arcsinx + alpha and its derivative is known to all:)

OpenStudy (anonymous):

no i havn't......i solve problems in my own style..not like just coping from wolframalpha or in a conventional way....thats my salient features...

OpenStudy (anonymous):

got it vaishali?

OpenStudy (anonymous):

we guys dont copy frm websites here...chill..calm down ppl

OpenStudy (anonymous):

I dont copy it from wolframa , but when someone does put up a really small answer to something that involves a ton of chain rules then I am a little suspect and I use wolframa to verify the correctness of the answer when I dont feel like doing all the mindless differentiation

OpenStudy (anonymous):

and wolframa does not agree with you at all

OpenStudy (anonymous):

For what it is worth, a Mathematica 8 solution with comments is attached.